solve by substitution method for 8m+n=17 and m-9n=66
Answer by Maths68(1474)

8m+n=17................(1)
m-9n=66................(2)
Take (1)
8m+n=17
n=17-8m................(3)
Substitute the value of n from (3) to (2)m-9n=66................(2)
m-9n=66................(2)
m-9(17-8m)=66
m-153+72m=66
73m=66+153
73m=219
73m/73=219/73
m=3
Plug in the value of m in (1)
8m+n=17................(1)
8(3)+n=17
24+n=17
n=17-24
n=-7

m,n=3,-7


Check
========
Plug in the value of m and n in (1)
8m+n=17................(1)
8(3)+(-7)=17
24-7=17
17=17

Plug in the value of m and n in (2)
m-9n=66................(2)
3-9(-7)=66
3+63=66
66=66